# Exploring the Physics of the Plasma Liner Experiment: A Multi-dimensional Study with FLASH, OSIRIS, and HELIOS

### v1

@@ IV Liner Formation
In the full experimental-scale PLX, the liner shell is formed using 36 plasma guns, but some of the key physics can be studied by simulating the merging of just two jets. The primary goal of our liner formation simulations was to understand which conditions lead to fluid-like regimes (collisional) and which lead to more kinetic-like regimes (collisionless). This jet-merging process in the context of PJMIF was recently studied by Cagas et al. (2023), and their results were consistent with some early findings from the PLX. Fluid codes (e.g., FLASH) will always produce shocks when the two jets collide, but these shocks can be mitigated through the collisionless interpenetration of the jets, a process that requires a kinetic model (i.e., not a fluid code). Here we have chosen to simulate the merging of two liner jets with both a kinetic code (OSIRIS) and a fluid code (FLASH) to make direct comparisons that help quantify the collisionality of the liner formation phase.
@@ IV.1 OSIRIS Simulations
A series of 1D OSIRIS simulations were conducted to quantify the interpenetration during the liner-liner jet merging process. A transition from a collisional to collisionless regime was observed as the initial jet velocity was increased. In these simulations, two singly-charged xenon plasma jets with density were initialized with equal electron and ion temperatures eV. The two jets counter-propagated with velocity , where is the controlled parameter. Fig. 3 shows the evolution of the plasma jet density, where the white lines correspond to the density profile of the left jet at the end of the simulation. At very high jet velocities (e.g., km/s), the merging of the two jets was collisionless; the left jet could fully interpenetrate through the right jet, as indicated by the white line in Fig. 3(a) (i.e., the final left jet density profile peaks on the right side of the domain). At very low jet velocities (e.g., km/s), the merging of the two jets was collisional; the left jet was mostly stopped by the right jet with minimal interpenetration as shown in Fig. 3(d). At intermediate velocities, the two jets can partially interpenetrate and form a broad density profile that peaks near the center, which is desirable for the PLX design. These results are consistent with the theoretical estimates of the ion mean free path, which is close to the liner jet size (5 cm) at km/s.
@@ IV.2 FLASH Simulations
In this section, 1D liner-liner jet merging simulations conducted with FLASH are presented and compared with those from OSIRIS. This comparison aims to better delineate the transition from collisional to collisionless regimes and to identify any additional hydrodynamic effects. The simulation setup for the two counter-propagating xenon jets is consistent with the previous subsection, aside from the introduction of a low-density pseudo-vacuum, with a density four orders of magnitude lower than the jet density. Both the jets and the pseudo-vacuum are modeled using an ideal gamma-law EOS, with . Thermal conduction and electron-ion heat exchange are described with Spitzer models. Spitzer (1962)
FLASH is able to simulate radiation transport using a multi-group diffusion model. The opacities for the jets and the pseudo-vacuum were tabulated using the PROPACEOS code. The initial radiation temperature was set to be in equilibrium with the ions and electrons, at eV. Simulations were conducted both with and without the inclusion of radiation transport.
Fig. 4 shows the evolution of total density in simulations using OSIRIS, FLASH without radiation transport, and FLASH with radiation transport, at four different jet velocities. In the collisional regime, observed at lower jet velocities ( 13 and 25 km/s), FLASH simulations without radiation transport show qualitative similarities with OSIRIS. However, the behavior diverges at higher velocities, with OSIRIS jets interpenetrating, while the FLASH jets compress to a higher density. Including radiation transport in FLASH modifies the dynamics, providing an additional energy reservoir for work done to compress the jets, leading to further collapse of the liner.
@@ V Target Compression
The kinetic process of target compression was investigated with a 1D OSIRIS simulation in Cartesian geometry. The initial condition of the PIC simulation is illustrated in Figs. 5(a) and 5(b). A 5 cm singly-charged () xenon liner propagated from right to left with velocity km/s and collided with a 20 cm stationary fully-ionized hydrogen fuel that was magnetized by a 50 T magnetic field. Reflecting boundary conditions were used for both fields and particles. The initial densities for the liner and fuel plasma were and , respectively. The simulation grid size was on the same order as the Debye length of the fuel with temperature eV, which consequently did not resolve the cold liner ( eV). To reduce the numerical heating of the liner, a high-order current-smoothing and field-smoothing scheme Shang (1999) was used. The total energy increased by about 7% at the peak compression of the fuel, which is much less than the change in fuel thermal energy. At , the majority of the energy was stored in the form of the liner kinetic energy. During the liner-fuel compression process, the liner kinetic energy was converted into thermal energy of the liner and fuel. Both the liner and fuel plasma were heated as evidenced by the broadening of the phase spaces along the axis shown in Figs. 5(c) and 5(d). As shown in Fig. 5(c), the magnetic field was also compressed by the liner, keeping the ratio nearly constant within the fuel region.
The evolution of the fuel density and temperature obtained from the PIC simulation agrees with the MHD simulations qualitatively as illustrated in Fig. 6. All three codes predicted peak compression to occur at about . The maximum temperature and density achieved in the OSIRIS simulation were slightly lower than those achieved in the FLASH and HELIOS simulations, which is likely due to mixing of the liner and fuel plasmas. This mixing also resulted in a shallower density gradient as shown in Fig. 6.
In the OSIRIS simulation, a low-pass filter was applied to the electromagnetic field in the region cm after to suppress a numerical instability observed in that region. Since the plasma density in that region is very low ( ) after , we do not anticipate the additional low-pass filter to influence the physics of the liner-fuel compression which happened in the region cm. Any meaningful influence should travel at the Alfven speed from cm to cm, which would take more than 1 . As shown in Fig. 6(a), the high temperature in the liner-fuel mixing region after the peak compression is likely an artifact due to an insufficient number of particles in that region.
@@
The Xe liner is initialized with a density of roughly g cm-3, temperature of 1.5 eV, velocity of 60 km/s, and thickness of 5 cm. For the simulation cases with liner perturbations, we imposed Gaussian density and velocity perturbations throughout the thickness of the liner at the angular positions of the experimental plasma guns. An example of this kind of perturbation is shown in Fig. 8, which displays the mathematical form (i.e., dimensionless) of the perturbation as a spherical slice at a radius of 25 cm (the inner liner radius) projected to a 2D map. The purpose of Fig. 8 is to illustrate that the regions of increased density and velocity are indeed centered at the angular positions of the experimental plasma guns, and that each Gaussian has an overall spot size larger than 4.25 cm (the initial jet radius at the chamber wall). These perturbations are normalized to have precise control of their amplitude, and we ran three simulations with 0%, 5%, and 10% perturbations.
Outside of the liner is a pseudo-vacuum, which we model as an extension of the Xe liner but with a density and velocity that decrease linearly to g cm-3 and 0 km/s, respectively. The initial conditions for these simulations are summarized in Table 1.
Parameter | Experimental-scale | Reactor-scale |
@@
We first look at the 0% perturbation case in detail to understand the dynamics and relevant physics. Fig. 9 shows a time series of the implosion, viewed as 2D slices of mass density (g cm-3) and ion temperature (eV). At the early time of 1.6 s, the target density has more than tripled from its initial value, and a temperature gradient has been established by the competing effects of compressional heating and thermal conduction. The central temperature has increased due to compression, but the target plasma’s contact with the colder liner plasma has resulted in some thermal losses. Magnetization helps reduce thermal losses to some extent, and ion temperatures overall increase. Target density and ion temperature both increase for the duration of the simulation; volume-averaged values reach 500 times the initial value and 150 eV, respectively. Localized peak values may be much higher, and it is important to keep in mind that results could be optimized on this platform and further improved on a larger reactor-scale platform with more plasma guns. Also note that, without any liner perturbations, the implosion remains spherical.
@@
Fig. 10 shows the same time series as the previous figure but now with magnetic field-relevant parameters: magnetic field magnitude (G), electron Hall parameter , and inverse plasma beta (the ratio of magnetic pressure to thermal pressure). There are a few key points to note here: 1) the randomization of the initial magnetic field results in peaks and valleys that make the plots appear noisy, 2) is overall less than unity as some B-field has been lost, and 3) remains very small during the simulation, i.e., the magnetic field is never dynamically important. Although has peak values 100, these are highly localized to where the magnetic field is also peaked, thus much of the overall benefit of magnetized thermal insulation is lost. In fact, the volume-averaged inside the target drops below unity at about 0.3 s. We can see an enhancement of B-field inside the liner at 3.4 s, which suggests that the field has diffused out of the target plasma in part because the cold liner has a relatively high magnetic resistivity.
@@
These results show that it is important for future PLX experiments to guard against magnetic field loss via resistive diffusion. In addition to simply using stronger initial fields, there are other ways this effect may be mitigated. More target preheat would decrease the target resistivity and help keep the field more “frozen-in” to the target plasma, allowing for more magnetic field compression. Keeping the liner hotter may also be beneficial, and could be achieved by using a material with a lower atomic number than Xe (). A lower- material would directly have a lower resistivity, but would also radiate less, keeping it hotter and thus also decreasing resistivity. Of course high- materials have their own benefits, namely their increased radiation losses make them colder and more effective pistons for compression. Thus there is some optimization to be done in this parameter space, but this is beyond the scope of this work.
@@
For these experimental-scale simulations, the minimum fuel radius that can be resolved with 10 computational cells is 1.25 cm, or a convergence ratio (CR) of 20. The target plasma in the simulations may continue to compress to higher CR, but results become less reliable due to the resolution. Also, higher CRs are not expected to be attainable experimentally, so we consider to be close to stagnation. Fig. 11 shows the mass density (g cm-3) and ion temperature (eV) for all perturbation cases at 4.0 s, which is when the 0% perturbation case reaches . We see that the 0% case is still mostly spherical at this late time, while the perturbed cases have become distorted. In the 5% case, the target plasma is still intact, but a significant portion of it has become mixed with colder liner material resulting in final target ion temperatures 2-3x less than the 0% case. In both the 0% and 5% cases, ion temperatures before stagnation reached peak values ranging from 150-200 eV, but the target remains hotter for longer in the 0% case. The target is completely destroyed in the 10% case as liner plasma has reached the origin. Based on these results, the 5% perturbation level is a reasonable upper limit on liner non-homogeneities for maintaining a viable target plasma. The uniformity of the liner will depend on the jet merging process and physical effects therein, some of which were discussed in Section IV.
@@
Fusion will not occur under the plasma conditions attained in these experimental-scale simulations, but that was expected. The current PLX platform is primarily for testing, optimizing, and exploring relevant physics. Our target compression simulation results show the importance of liner uniformity and that ion temperatures of about 200 eV should be possible, and perhaps higher temperatures can be reached with further optimization. However, it has always been the case that PLX would need to be scaled to a larger system to be a viable option for fusion energy production, which was the impetus for the simulation presented in the next subsection.
@@ V.3 Reactor-scale FLASH Simulation
Here we present a 3D FLASH target compression simulation using idealized initial conditions, which are summarized in Table 1 in the previous subsection. The chosen parameters represent theoretically attainable conditions for a larger, reactor-scale PLX design that employs many more plasma guns for both target and liner formation, hence the much higher densities as compared to the previous experimental-scale simulations. Furthermore, the reactor-scale simulation assumes that some perfectly-spherical compression has already taken place; the initial target radius is only 4 cm, and the initial liner thickness is 1 cm. This was done, in part, to accurately model the problem with higher resolution (a minimum cell width of 0.03125 cm) and track the compression down to smaller radii, or higher convergence ratios. If we consider 25 cm to still be the “true” initial target radius, then this simulation is starting at . We should also note that the target plasma in this simulation is 50/50 deuterium-tritium (DT) as opposed to pure hydrogen, and it is initialized with a higher temperature of 108 eV.
@@
The initial magnetic field in the target is randomized with the same method as described in the previous subsection. We purposely aimed for a regime in which and were both close to 10, which is similar to the experimental-scale regime. Due to the increased density and temperature, the resulting volume-averaged field strength required to reach this regime was roughly 6.29 T.
@@
As with the experimental-scale simulations of the previous subsection, the results of the reactor-scale simulation are reliably accurate as long as the target radius is resolved with at least 10 computational cells. This means we can resolve a minimum fuel radius of 0.3125 cm, which corresponds to an overall . We consider this CR to be close to stagnation, and it is reached at roughly 320 ns. Fig. 12 shows the mass density (g cm-3) and ion temperature (eV) at this time, and we observe that the target plasma has reached fusion-relevant conditions. Peak ion temperatures in the target have exceeded 1 keV, and the volume-averaged ion number density has increased by nearly a factor of 500. It is also worth noting that the implosion has remained roughly spherical to this high CR, which is perhaps not too surprising since liner perturbations were not used.
@@
Like the experimental-scale simulations, the target plasma in the reactor-scale simulation lost some magnetic field due to resistive diffusion, but the effect was less dramatic. Despite starting with a lower electron Hall parameter, as compared to the experimental-scale case, does not go below unity until about 247 ns, which corresponds to a . This is much later in the compression compared to the experimental-scale 0% perturbation case, which reached this condition at a CR slightly greater than 1.0. There are multiple reasons for this including: 1) the reactor-scale liner is moving faster, giving less time for B-field to be lost, and 2) the reactor-scale target is hotter, which results in a lower magnetic resistivity. This supports the idea, presented in the previous subsection, that some magnetic field losses can be mitigated by increasing the target preheat during target formation. Using a faster liner may benefit the magnetic field, but too much interpenetration of the liner jets may not be desirable for PLX as mentioned in Section IV.1.

### v2

@@ IV Liner Formation
In the full experiment-scale PLX, the liner shell is formed using 36 plasma guns, but some of the key physics can be studied by simulating the merging of just two jets. The primary goal of our liner formation simulations was to understand which conditions lead to fluid-like, collisional regimes and which lead to more kinetic-like, collisionless regimes. This jet-merging process in the context of PJMIF was recently studied by Cagas2023, and their results were consistent with some early findings from the PLX. Fluid codes (such as FLASH) will always produce shocks when the two jets collide, but these shocks can be mitigated through the collisionless interpenetration of the jets, a process that requires kinetic modeling.
Here we have chosen to simulate the merging of two liner jets with both a particle-in-cell code (OSIRIS) and a fluid code (FLASH) to make direct comparisons that help quantify the collisionality of the liner formation phase.
@@ IV.1 OSIRIS Simulations
A series of 1D OSIRIS simulations were conducted to quantify the interpenetration during the liner-liner jet merging process. A transition from a collisional to collisionless regime was observed as the initial jet velocity was increased. In these simulations, two singly-charged xenon plasma jets with density were initialized with equal electron and ion temperatures eV. The two jets counter-propagated with velocity , where is the controlled parameter. Fig. 3 shows the evolution of the plasma jet density, where the white lines correspond to the density profile of the left jet at the end of the simulation. At high jet velocities (e.g., cm ), the merging of the two jets was collisionless; the left jet was found to fully penetrate through the right jet, as indicated by the white line in Fig. 3(a) (i.e., the final left jet density profile peaks on the right side of the domain). At low jet velocities (e.g., cm ), the merging of the two jets was collisional; the left jet was mostly stopped by the right jet with minimal interpenetration as shown in Fig. 3(d). At intermediate velocities, the two jets can partially interpenetrate and form a broad density profile that peaks near the center, which is desirable for the PLX design. These results are consistent with the theoretical estimates of the ion mean free path, which is close to the liner jet size (5 cm) at cm .
@@ IV.2 FLASH Simulations
In this section, 1D liner-liner jet merging simulations conducted with FLASH are presented and compared with those from OSIRIS. This comparison aims to better delineate the transition from collisional to collisionless regimes and to identify any additional effects from radiative losses. The simulation setup for the two counter-propagating xenon jets is consistent with the previous subsection, aside from the introduction of a low-density pseudo-vacuum, with a density four orders of magnitude lower than the jet density. Both the jets and the pseudo-vacuum are modeled using an ideal gamma-law EOS, with . Thermal conduction and electron-ion heat exchange are described with Spitzer models. Spitzer1962
To incorporate radiation effects, we use FLASH’s multi-group diffusion radiation transport model. The opacities for the jets and the pseudo-vacuum were tabulated using the PROPACEOS code. The initial radiation temperature was set to be in equilibrium between ions and electrons, at eV. Simulations were conducted both with and without the inclusion of radiation transport to assess the impact of the latter.
Fig. 4 shows the evolution of total density in simulations using OSIRIS, FLASH without radiation transport, and FLASH with radiation transport, at four different jet velocities. In the collisional regime, observed at lower jet velocities ( and cm ), FLASH simulations without radiation transport show qualitative similarities with the OSIRIS results. However, the behavior diverges at higher velocities, with OSIRIS jets interpenetrating while the FLASH jets compressing to a higher density. Including radiation transport in FLASH modifies the dynamics of the interaction: the addition of radiative losses allows for more compression, leading to further collapse of the liner. This indicates that while the particle-in-cell modeling can more accurately capture the regime transition from collisional to collisionless, radiation losses may also be important during the liner formation.
@@ V Target Compression
The kinetic process of target compression was investigated with a 1D OSIRIS simulation in Cartesian geometry. The initial condition of the PIC simulation is illustrated in Figs. 5(a) and 5(b). A 5 cm singly-charged () xenon liner propagated from right to left with velocity and collided with a 20-cm, stationary, fully-ionized hydrogen fuel that was magnetized by a 50 T magnetic field. Reflecting boundary conditions were used for both fields and particles. The initial densities for the liner and fuel plasma were and , respectively. The simulation grid size was on the same order as the Debye length of the fuel with temperature eV, which consequently did not resolve the cold liner ( eV). To reduce the numerical heating of the liner, a high-order current-smoothing and field-smoothing scheme Shang1999 was used. The total energy increased by about 7% at the peak compression of the fuel, which is much less than the change in fuel thermal energy. At , the majority of the energy was stored in the form of the liner kinetic energy. During the liner-fuel compression process, the liner kinetic energy was converted into thermal energy of the liner and fuel. Both the liner and fuel plasma were heated as evidenced by the broadening of the phase spaces along the axis shown in Figs. 5(c) and 5(d). As shown in Fig. 5(c), the magnetic field was also compressed by the liner, keeping the ratio nearly constant within the fuel region.
The evolution of the fuel density and temperature obtained from the PIC simulation qualitatively agrees with the MHD simulations as illustrated in Fig. 6. All three codes predicted peak compression to occur at about . The maximum temperature and density achieved in the OSIRIS simulation were slightly lower than those achieved in the FLASH and HELIOS simulations, which is likely due to mixing of the liner and fuel plasmas. This mixing also resulted in a shallower density gradient as shown in Fig. 6.
In the OSIRIS simulation, a low-pass filter was applied to the electromagnetic field in the region cm after to suppress a numerical instability observed in that region. Since the plasma density in that region is very low ( ) after , we do not anticipate the additional low-pass filter to influence the physics of the liner-fuel compression which happened in the region cm. Any meaningful influence should travel at the Alfvèn speed from cm to cm, which would take more than 1 . As shown in Fig. 6(a), the high temperature in the liner-fuel mixing region after the peak compression is likely an artifact due to an insufficient number of particles in that region.
@@
The Xe liner is initialized with a density of roughly g cm-3, temperature of 1.5 eV, velocity of cm , and thickness of 5 cm. For the simulation cases with liner perturbations, we imposed Gaussian density and velocity perturbations throughout the thickness of the liner at the angular positions of the experimental plasma guns. An example of this kind of perturbation is shown in Fig. 8, which displays the mathematical form (dimensionless) of the perturbation as a spherical slice at a radius of 25 cm (the inner liner radius), projected to a 2D map. The purpose of Fig. 8 is to illustrate that the regions of increased density and velocity are indeed centered at the angular positions of the experimental plasma guns, and that each Gaussian has an overall spot size larger than 4.25 cm (the initial jet radius at the chamber wall). These perturbations are normalized to have precise control of their amplitude. Three simulations were executed with 0%, 5%, and 10% perturbations.
Outside of the liner is a pseudo-vacuum, which we model as an extension of the Xe liner but with a density and velocity that decrease linearly to g cm-3 and 0 cm , respectively. The initial conditions for these simulations are summarized in Table 1.
| Parameter | Experiment-scale | Reactor-scale |
@@
We first look at the 0% perturbation case in detail to understand the dynamics and relevant physics. Fig. 9 shows a time series of the implosion, viewed as 2D slices of mass density (g cm-3) and ion temperature (eV). At the early time of 1.6 s, the target density has more than tripled from its initial value, and a temperature gradient has been established by the competing effects of compressional heating and thermal conduction. The central temperature has increased due to compression, but the target plasma’s contact with the colder liner plasma has resulted in some thermal losses. Magnetization helps reduce thermal losses to some extent, and ion temperatures overall increase. Target density and ion temperature both increase for the duration of the simulation; volume-averaged values reach 500 times the initial value and 150 eV, respectively. Localized peak values can be higher and the platform can be further optimized to improve performance on a larger reactor-scale platform with more plasma guns. Without any liner perturbation, the implosion remains spherical.
@@
Fig. 10 shows the same time series as the previous figure but now for magnetic field relevant parameters: magnetic field magnitude (G), electron Hall parameter , and inverse plasma beta (the ratio of magnetic-to-thermal pressure). There are a few key points of note: 1) the randomization of the initial magnetic field results in peaks and valleys that make the plots appear noisy; 2) is overall less than unity as some B-field has been lost; and 3) remains very small during the simulation, i.e., the magnetic field is never dynamically important. Although has peak values 100, these are highly localized to where the magnetic field is also peaked, thus much of the overall benefit of magnetized thermal insulation is lost. In fact, the volume-averaged inside the target drops below unity at about 0.3 s. We can see an enhancement of B-field inside the liner at 3.4 s, which suggests that the field has diffused out of the target plasma in part because the cold liner has a relatively high magnetic resistivity.
@@
These results suggest that future PLX experiments should guard against magnetic field loss due to resistive diffusion. In addition to simply using stronger initial fields, there are other ways this effect may be mitigated. More target preheat would decrease the target resistivity and help keep the field more “frozen-in” to the target plasma, allowing for more magnetic field compression. Keeping the liner hotter may also be beneficial, and could be achieved by using a material with a lower atomic number than Xe. A lower- material would inherently have a lower resistivity, but would also radiate less, maintain a relatively smaller resistivity. These effects must be balanced against the benefits of the cold-liner performance as an effective piston for for compression. There is some room for optimization in this parameter space, but such a task is beyond the scope of this work.
@@
For these experiment-scale simulations, the minimum fuel radius that can be resolved with 10 computational cells is 1.25 cm, or a convergence ratio (CR) of 20. The target plasma in the simulations may continue to compress to higher CR, but results become less reliable due to resolution constraints. Also, higher CRs are not expected to be attainable experimentally, so we consider to be close to stagnation. Fig. 11 shows the mass density (g cm-3) and ion temperature (eV) for all perturbation cases at 4.0 s, which is when the 0% perturbation case reaches . We see that the 0% case is still mostly spherical at this late time, while the perturbed cases have become distorted. In the 5% case, the target plasma is still present but a significant portion has become mixed with the colder liner material resulting in final target ion temperatures below the 0% case. In both the 0% and 5% cases, ion temperatures before stagnation reached peak values ranging from 150-200 eV, but the target remains hotter for longer in the 0% case. The target is found to be completely disrupted in the 10% case as liner plasma has reached the origin. Based on these results, the 5% perturbation level is a reasonable upper limit on liner non-homogeneities for maintaining a viable target plasma. The uniformity of the liner will depend on the jet merging process and physical effects therein, some of which were discussed in Section IV.
@@
Fusion will not occur under the plasma conditions attained in these experiment-scale simulations, but that was expected. The current instance of the PLX platform is primarily for testing, optimizing, and exploring relevant physics. Our target compression simulation results show the importance of liner uniformity and that ion temperatures of 200 eV should be possible. Higher temperatures may be within reach with further optimization. To be a viable option for fusion energy production, the PLX would need to be scaled to a larger system.
@@ V.3 Reactor-scale FLASH Simulation
Here we present a 3D FLASH target compression simulation using idealized initial conditions, which are summarized in Table 1. The chosen parameters represent theoretically attainable conditions for a larger, reactor-scale PLX design that employs many more plasma guns for both target and liner formation, hence the much higher densities as compared to the previous experiment-scale simulations. Furthermore, the reactor-scale simulation assumes that some perfectly-spherical compression has already taken place; the initial target radius is only 4 cm, and the initial liner thickness is 1 cm. This was done, in part, to accurately model the problem with higher resolution (a minimum cell width of 0.03125 cm) and track the compression down to smaller radii, or higher convergence ratios. If we consider 25 cm to still be the “true” initial target radius, then this simulation is starting at . We should also note that the target plasma in this simulation is 50/50 deuterium-tritium (DT) as opposed to pure hydrogen, and it is initialized with a higher temperature of 108 eV.
@@
The initial magnetic field in the target is randomized with the same method as described in the previous subsection. We purposely aimed for a regime in which and were both close to 10, which is similar to the experiment-scale regime. Due to the increased density and temperature, the resulting volume-averaged field strength required to reach this regime was roughly 63 kG.
@@
As with the experiment-scale simulations of the previous section, the results of the reactor-scale simulation are reliably accurate as long as the target radius is resolved with at least 10 computational cells. This means we can resolve a minimum fuel radius of 0.3125 cm, which corresponds to an overall . We consider this CR to be close to stagnation, and it is reached at roughly 320 ns. Fig. 12 shows the mass density (g cm-3) and ion temperature (eV) at this time, and we observe that the target plasma has reached fusion-relevant conditions. Peak ion temperatures in the target have exceeded 1 keV, and the volume-averaged ion number density has increased by nearly a factor of 500. It is also worth noting that the implosion has remained roughly spherical to this high CR, which is perhaps not too surprising since the liner was not perturbed.
@@
Like in the experiment-scale simulations, the target plasma in the reactor-scale simulation lost some magnetic field due to resistive diffusion, but the effect was less pronounced. Despite starting with a lower electron Hall parameter, as compared to the experiment-scale case, does not go below unity until about 247 ns, which corresponds to a . This is much later in the compression compared to the experiment-scale 0% perturbation case, which reached this condition at a CR slightly greater than 1.0. This is because 1) the reactor-scale liner is moving faster, giving less time for B-field to be lost, and 2) the reactor-scale target is hotter, which results in lower magnetic resistivity. This supports the idea, presented in the previous section, that some magnetic field losses can be mitigated by increasing the target preheat during target formation. Using a faster liner may also benefit the magnetic field, but too much interpenetration of the liner jets may not be desirable for the PLX, as mentioned in Section IV.1.